# Agentless [Uncaptioned image] : Demystifying LLM-based Software Engineering Agents

### v1

@@ 7 Conclusion
We propose Agentless– an agentless approach to automatically tackle software development problems. Agentless uses a simple two phase approach of localization followed by repair. Compared to prior agent-based approaches, Agentless deliberately disallows the LLM for autonomous tool usage or planning. Our evaluation on the popular SWE-bench Lite benchmark demonstrates that Agentless can achieve the highest performance compared with other open-source techniques while at the same time minimizing the cost. Furthermore, we perform a detailed classification of problems in SWE-bench Lite to not only offer new insights but to construct a more rigorous benchmark of SWE-bench Lite-$S$ after removing problematic problems.
@@ Acknowledgments
We thank Jiawei Liu for providing some of the resources used to run the experiments. One of the authors would like to thank Jun Yang for generously gifting his old bike [^3] Sadly the bike is currently broken. which allowed the author to travel faster and thus increasing research speed.

### v2

@@ 8 Conclusion
We propose Agentless– an agentless approach to automatically tackle software development problems. Agentless uses a simple three phase approach of localization, repair, and patch validation. Compared to prior agent-based approaches, Agentless deliberately disallows the LLM for autonomous tool usage or planning. Our evaluation on the popular SWE-bench Lite benchmark demonstrates that Agentless can achieve the highest performance compared with other open-source techniques while at the same time minimizing the cost. Furthermore, we perform a detailed classification of problems in SWE-bench Lite to not only offer new insights but to construct a more rigorous benchmark of SWE-bench Lite-$S$ after removing problematic problems.
@@ Acknowledgments
We thank Jiawei Liu and Yuxiang Wei for providing some of the resources used to run the experiments. One of the authors would like to thank Jun Yang for generously gifting his old bike [^4] Sadly the bike is currently broken. which allowed the author to travel faster and thus increasing research speed.